# Tracking Protostellar Variability in Massive Protoclusters with ALMA: I. Insights from QUARKS and MaMMOtH

### v1

@@ 4.3.2 Difference Maps Analysis
Here we present the results for I13111–6228, the most significant case in our sample. A representative example is shown in Appendix D, while the complete set of figures is available online. Panels (a) and (b) of Figure 7 show the 1.3 mm continuum emission toward I13111–6228 observed with ALMA in May 2023 and June 2024. The two datasets were processed with identical calibration and imaging procedures to ensure a fair comparison. Panel (c) presents the difference map formed by subtracting panel (a) from panel (b). A compact residual feature is detected at the position of condensation 10, while no significant residuals appear toward the other condensations. The peak intensity of condensation 10 increases from in 2023 to in 2024 after relative calibration, corresponding to a 68% rise with an uncertainty of 2%. The rms noise level of the difference map is , and the residual peak reaches (S/N 48). The large, isolated residual confirms that condensation 10 in I13111–6228 is a genuinely variable source.
To examine the physical nature of the candidate variables, we used molecular line emission from four spectral windows in the cleanest dataset from the QUARKS survey, combining continuum data from C-5 (TM1, ″), C-2 (TM2, ″), and ACA 7-m array ( ″). Condensations that exhibit H30 emission were classified as Hii regions, while those without H30 but exhibiting CH3CN emission were classified as hot cores. Among the Hii regions, those with extremely compact sizes ( pc) and high electron densities ( cm-3) are further categorized as hypercompact (HC) Hii regions (2005Kurtz). In addition to their physical compactness, HC Hii regions also differ from ultracompact (UC) Hii regions in that they typically have broader radio recombination line (RRL) widths, ranging from 40–100 km s-1, instead of the typical 25–30 km s-1 found in most UC Hii regions (2019Yang).
Based on this classification scheme, condensation 10 in I13111–6228 is identified as a HC Hii region. Figure 8 shows the integrated intensity (moment 0) map of the H30 emission. This source displays a broad H30 linewidth (40 km s-1). Owing to its compact size (0.01 pc) and the large velocity width of its recombination line, we classify it as a HC Hii region. The other four candidates are classified as hot cores. The full combination method and classification strategy will be presented in Jiao et al. (in prep.). The physical properties of these variables will be examined in detail in forthcoming studies.
5 Discussion
@@
As described in Sect. 4.3, we identified five variable sources among the 383 condensations using a conservative variability metric, with one source standing out as the most significant case. This corresponds to a variable detection rate of 1.3%. The identification is based on standard deviation analysis and is further confirmed through continuum image difference maps. In the following, we briefly discuss the implications of these findings and our future plan.
Figure 9 presents a scatter plot of the fractional amplitude as a function of the mean peak intensity for all sources with valid multi-epoch millimeter measurements. The background histogram displays the distribution of mean peak intensities across the full sample, providing context for the population’s brightness. Each gray point denotes an individual condensation, while the red circles highlight the detected variables. The temporal sampling of only two to four epochs is insufficient to trace long-term variability. Detections are also biased toward brighter condensations due to signal-to-noise constraints, suggesting that faint sources require larger intrinsic variations to exceed our detection threshold. As a result, the measured variability detection rate only gives a lower limit of the real value.
@@
The ability to detect (sub)millimeter continuum variability in massive protoclusters is also fundamentally limited by sensitivity and angular resolution. Single-dish telescopes (e.g., JCMT, IRAM 30-m) have beam sizes of several to tens of arcseconds, encompassing multiple protostellar cores and diffuse emission within a single field of view. Consequently, intrinsic variations from individual sources are spatially averaged, leading to severe beam dilution. Previous JCMT studies have emphasized this limitation, showing that high-resolution and high-sensitivity observations can detect flux variability and identify the embedded young stellar objects responsible for such events more effectively (2019Park).
@@
In contrast, high-resolution interferometric observations are able to resolve individual protostars and measure their flux variations in protoclusters at large distances. Therefore, building on our methodology and findings, the next step is to use the growing ALMA multi-epoch archive for a larger variability search. Our efforts will proceed along two main paths. First, we plan to expand the variability sample using the ALMA archive. The ALMA Evolutionary Study of High-Mass Protocluster Formation in the Galaxy (ALMAGAL; PIs: Sergio Molinari, Peter Schilke, Cara Battersby, Paul Ho; Project ID: 2019.1.00195.L; 2025Molinari) survey shares band 6 similar array configurations with QUARKS, specifically the 12-m C–5 configuration (TM1, 0.3″). By cross-matching the two datasets, we find 57 overlapping clumps, providing an opportunity to increase the statistical significance of the variability sample. In addition to enlarging the sample, we plan to incorporate higher-resolution data from surveys such as the Digging into the Interior of Hot Cores with ALMA (DIHCA; PI: Sanhueza; Project ID: 2016.1.01036.S) survey to investigate substructures and drivers of physical variability on smaller spatial scales. Second, we aim to obtain new multi-epoch observations to expand temporal coverage and to follow up on the candidate variable sources identified in this work. These combined efforts will lay a solid foundation for constraining the occurrence rate, amplitude, and physical mechanisms of millimeter variability during the early phases of massive proto-cluster evolution.
6 Conclusion
@@ Acknowledgements
T.L. acknowledges the supports by the National Key R&D Program of China (no. 2022YFA1603100), National Natural Science Foundation of China (NSFC) through grants no. 12073061 and no. 12122307, and the Tianchi Talent Program of Xinjiang Uygur Autonomous Region.
@@ Appendix A Observing Parameters
| ID | Source NameaaCorresponds to the “Condensation ID” entries listed in Table B1. | R.A. (J2000) | Decl. (J2000) | Calibrators | Min./Max. BL | Obs. Date | Beam SizebbCorresponds to the “Obs. Date” entries listed in Table A1. | |
| (h:m:s) | (d:m:s) | Phase | Bandpass/Flux | (m/m) | (yyyy-mm-dd) | () | ||
| 1 | G294.521.62 | 11:35:34.06 | :14:49.30 | J11236417 | J11074449 | 15.1/783.5 | 2022-06-02 | |
| 15.3/1210.6 | 2022-08-21 | |||||||
| I113326258 | 11:35:32.23 | :14:46.80 | J10476217 | J11074449 | 15.1/1397.8 | 2024-05-31 | ||
| 2 | I123206122 | 12:34:53.38 | :39:46.90 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 3 | I123266245 | 12:35:34.81 | :02:32.10 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 4 | I123836128 | 12:41:17.32 | :44:38.60 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 5 | I1257263161 | 13:00:24.03 | :32:31.90 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 6 | I1257263162 | 13:00:28.73 | :32:37.30 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 7 | G305.21+0.21 | 13:11:13.78 | :34:41.90 | J13376509 | J14274206 | 15.1/783.5 | 2022-06-02 | |
| I1307962181 | 13:11:13.73 | :34:40.20 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | ||
| 15.1/1397.8 | 2024-06-02 | |||||||
| 8 | I1307962182 | 13:11:09.50 | :34:39.70 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 9 | I130806229 | 13:11:14.28 | :44:58.30 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 10 | I131116228 | 13:14:26.49 | :44:28.30 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 11 | I131346242 | 13:16:42.99 | :58:29.30 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 12 | I131406226 | 13:17:15.90 | :42:27.00 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 13 | I1329162291 | 13:32:31.77 | :45:11.80 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 14 | I1329162292 | 13:32:34.58 | :45:27.00 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 15 | I132916249 | 13:32:31.23 | :05:21.80 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 16 | I132956152 | 13:32:53.49 | :07:49.30 | J13376509 | J16175848 | 15.3/2516.8 | 2023-05-02 | |
| 15.1/1397.8 | 2024-06-02 | |||||||
| 17 | G309.92+0.48 | 13:50:41.91 | :35:10.20 | J13376509 | J14274206 | 15.1/783.5 | 2022-06-02 | |
| I134716120 | 13:50:42.10 | :35:14.90 | J14085712 | J16175848 | 15.1/1397.8 | 2024-06-04 | ||
| 18 | G328.81+0.63 | 15:55:48.71 | :43:06.40 | J16034904 | J16175848 | 15.1/783.5 | 2022-05-30 | |
| 15.1/1301.6 | 2022-08-09 | |||||||
| I155205234 | 15:55:48.39 | :43:09.80 | J16034904 | J16175848 | 15.1/1397.8 | 2024-06-02 | ||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 19 | G331.130.24 | 16:10:59.68 | :50:15.50 | J16034904 | J16175848 | 15.1/783.5 | 2022-05-30 | |
| 15.1/1301.6 | 2022-08-09 | |||||||
| I160715142 | 16:10:59.01 | :50:21.60 | J16034904 | J16175848 | 15.1/1397.8 | 2024-06-02 | ||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 20 | G331.280.19 | 16:11:26.48 | :41:56.90 | J16034904 | J16175848 | 15.1/783.5 | 2022-05-30 | |
| 15.1/1301.6 | 2022-08-09 | |||||||
| I160765134 | 16:11:27.2 | :41:56.90 | J16034904 | J16175848 | 15.1/1397.8 | 2024-06-02 | ||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 21 | G336.990.03 | 16:35:33.43 | :31:11.60 | J16314345 | J16175848 | 15.1/783.5 | 2022-05-30 | |
| I163184724 | 16:35:33.20 | :31:11.30 | J16505044 | J16175848 | 15.1/1397.8 | 2024-06-02 | ||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 22 | G339.881.26 | 16:52:04.83 | :08:34.40 | J16314345 | J16175848 | 15.1/783.5 | 2022-05-30 | |
| I164844603 | 16:52:03.99 | :08:24.60 | J16505044 | J16175848 | 15.1/1397.8 | 2024-06-02 | ||
| 15.1/1397.8 | 2024-06-03 | |||||||
| 15.1/1397.8 | 2024-06-03 | |||||||
Appendix B Peak Intensities

### v3

@@ IV.3.2 Difference Maps Analysis
Here we present the results for I13111–6228, the most significant case in our sample. The difference maps for the remaining four variable sources are shown in Appendix E. Panels (a) and (b) of Figure 7 show the 1.3 mm continuum emission toward I13111–6228 observed with ALMA in May 2023 and June 2024. The two datasets were processed with identical calibration and imaging procedures to ensure a fair comparison. Panel (c) presents the difference map formed by subtracting panel (a) from panel (b). A compact residual feature is detected at the position of condensation 10, while no significant residuals appear toward the other condensations. The peak intensity of condensation 10 increases from in 2023 to in 2024 after relative calibration, corresponding to a 68% rise with an uncertainty of 2%. The rms noise level of the difference map is , and the residual peak reaches (S/N 48). The large, isolated residual confirms that condensation 10 in I13111–6228 is a genuinely variable source.
V Discussion
@@
As detailed in Sect. IV.3, we identified five variable sources from the total sample of 383 condensations by applying a conservative variability metric to their peak intensities. Among these, I13111–6228 stands out, exhibiting a 68% increase in continuum peak intensity over one year, with an uncertainty of 2%. The identification is based on a standard deviation analysis and is further confirmed through continuum image difference maps, from which the variability amplitudes are quantified. In the following, we briefly discuss the implications of these findings and outline directions for future work.
V.1 Detection Fraction and Sampling Biases
In this work, we derive a millimeter continuum variability detection fraction of 1.3% within our ALMA sample. This value should be interpreted as a lower limit on the intrinsic incidence of continuum variability, as it is strongly dependent on observational factors such as temporal sampling, sensitivity, angular resolution, and sample selection.
The sample analyzed in this work is drawn from two surveys with distinct target selection strategies. The QUARKS survey was designed to investigate star formation in a broadly unbiased sample of infrared-bright massive protoclusters that mostly host UC Hii region candidates, whereas the MaMMOtH survey specifically targets regions associated with 6.7 GHz methanol masers to study millimeter continuum variability. Consequently, the sample considered in this work is biased toward massive star-forming systems that are more likely to host actively accreting protostars or UC Hii regions.
The detection fraction is measured from 383 compact condensations identified across 22 massive protoclusters observed with ALMA Band 6 (1.3 mm) at an angular resolution of . Fourteen protoclusters were observed by both QUARKS and MaMMOtH, with eight additional regions observed exclusively by QUARKS. The protoclusters span molecular gas masses of 68–7585 and distances from 1.4 to 11.6 kpc (Liu et al., 2024; Xu et al., 2024); see Table A1.
All sources have at least two observing epochs separated by more than one year. Temporal sampling represents a further limitation: most condensations are observed in only two epochs, and at most four, which is insufficient to trace long-term variability. As a result, the intrinsic timescales of continuum variability cannot be robustly constrained with the current data.
Figure 8 presents the fractional amplitude as a function of the mean peak intensity for all condensations with valid multi-epoch millimeter measurements. Each point represents a single condensation, where the mean peak intensity is calculated by averaging the peak intensities over all available observing epochs. The background histogram shows the distribution of this source-averaged mean peak intensity across the full sample, providing context for the underlying brightness distribution. All five detected variable sources have source-averaged mean peak intensities exceeding 15 mJy beam-1 and are located among the brightest condensations in our sample.
The JCMT Transient Survey provides a useful point of reference for assessing the impact of observational strategy on variability detection. It has monitored submillimeter continuum variability since 2015 with a typical monthly cadence. Focusing primarily on nearby (500 pc) Gould Belt clouds dominated by low-mass star formation, the survey reports that more than 30% of bright sub-millimeter continuum sources (peak brightness 0.5 Jy beam-1 with a beam size of 14.6″) exhibit significant secular variability at 850 m (Lee et al., 2021).
Overall, the observed 1.3% detection fraction primarily reflects observational limitations and sample selection effects. It should therefore be regarded as an observationally defined lower limit on the incidence of millimeter continuum variability under the conditions of this study. Therefore, long-term monitoring observations as the JCMT Transient Survey did are needed to more quantitatively evaluate the variability of protostars within massive proto-clusters.
V.2 Classification of Variable Sources
To examine the physical nature of the detected variable sources, we used molecular line emission from four spectral windows in the cleanest dataset from the QUARKS survey, combining continuum data from C-5 (TM1, ″), C-2 (TM2, ″), and ACA 7-m array ( ″). Condensations exhibiting H30 emission are classified as Hii regions, while those without detectable H30 but showing strong CH3CN emission are classified as hot cores. Hypercompact (HC) Hii regions are characterized by extremely compact sizes ( pc) and high electron densities ( cm-3; Kurtz 2005). They also typically exhibit broader radio recombination line (RRL) FWHMs (40–100 km s-1) than ultracompact (UC) Hii regions, which usually show linewidths of 25–30 km s-1 (Yang et al., 2019).
Among the five variable sources, the most significant variability is observed in condensation 10 within I13111–6228, which is located in the G305 massive star-forming complex at a distance of kpc (Liu et al., 2024). It has a molecular gas reservoir of on clump scale (radius pc; Liu et al. 2024; Xu et al. 2024). Figure 9 presents the integrated intensity (moment 0) map of the H30 emission, with the corresponding spectrum shown in the inset. Given its compact physical size ( pc) and broad recombination line ( km s-1), we classify this source as a (HC) Hii region. We note that radio stars (e.g., MWC 349) can also exhibit broad recombination lines and significant variability (Gordon et al., 2001), our target could also be one of its kind. Although the exact nature of this source remains intriguing and requires further observations for confirmation, we tentatively classify it as a (HC) Hii region in this study.
Figure 10 shows the spectra for the five variable sources extracted from the four spectral windows of the QUARKS dataset. CH3CN emission is detected toward the other four variable sources. Among them, condensation 7 in I11332–6258, condensation 6 in I16076–5134, and condensation 1 in I16484–4603 exhibit rich complex molecular line emission and are thus classified as hot molecular cores.
Although CH3CN line emission is also detected toward condensation 5 in I12320–6122, its spectrum is significantly less line-rich compared to the other three hot molecular cores. As shown in Figure 11, this source, however, drives a highly energetic and collimated bipolar outflow. Based on this evidence, we classify it as a high-mass protostellar object (HMPO), which represents an earlier evolutionary phase than the hot core stage. In addition to I12320–6122, both I11332–6258 condensation 7 and I16484–4603 condensation 1 are also associated with bipolar molecular outflows (Jiao and others, in prep.). While condensation 6 in I16076–5134 is associated with a candidate explosive outflow (Guzmán Ccolque et al., 2022). The physical properties and variability mechanisms of these sources will be examined in further detail in our forthcoming studies.
V.3 Crowding Effect and the Need for Interferometric Monitoring
@@
The ability to detect (sub)millimeter continuum variability in massive protoclusters is also fundamentally limited by sensitivity and angular resolution. Single-dish telescopes (e.g., JCMT, IRAM 30-m) have beam sizes of several to tens of arcseconds, encompassing multiple protostellar cores and diffuse emission within a single field of view. Consequently, intrinsic variations from individual sources are spatially averaged, leading to severe beam dilution. Previous JCMT studies have emphasized this limitation, showing that high-resolution and high-sensitivity observations can detect flux variability and identify the embedded young stellar objects responsible for such events more effectively (Park et al., 2019).
@@
In contrast, high-resolution interferometric observations are able to resolve individual protostars and measure their flux variations in protoclusters at large distances. Therefore, building on our methodology and findings, the next step is to use the growing ALMA multi-epoch archive for a larger variability search.
VI Conclusion
@@ Acknowledgements
T.L. acknowledges the supports by the National Key R&D Program of China no. 2022YFA1603100, National Science and Technology Major Project 2024ZD1100601, National Natural Science Foundation of China (NSFC) through grants no. 12073061 and no. 12122307, the Tianchi Talent Program of Xinjiang Uygur Autonomous Region and the Tianshan Talent Training Program 2024TSYCTD0013.
@@
G.G. acknowledges support from the ANID Basal project FB210003.
S.Z. acknowledges support from the NAOJ ALMA Scientific Research Grant Code 2025-29B.
D.J. is supported by NRC Canada and by an NSERC Discovery Grant.
W.J. acknowledges support from the Shanghai Post-doctoral Excellence Program.
L.B.acknowledges support from the ANID Basal project FB210003.
DAL and SYP were supported by the Ministry of Science and Higher Education of the Russian Federation (theme No. FEUZ-2026-0012).
This work has been supported by the grant PID2024-155316NB-I00 funded by MICIU /AEI /10.13039/501100011033 / FEDER, UE and CSIC PIE 202350E189. This work was also supported by the Spanish program Unidad de Excelencia María de Maeztu financed by MCIN/AEI/10.13039/501100011033, and by the MaX-CSIC Excellence Award MaX4-SOMMA-ICE.
@@
We are grateful to the anonymous referee for valuable comments that significantly improved the quality of this paper.
@@ Appendix A Observing Parameters
| ID | Source NameaaCorresponds to the “Condensation ID” entries listed in Table C1. | R.A. (J2000) | Decl. (J2000) | Calibrators | Min./Max. BL | Obs. Date | Beam SizebbCorresponds to the “Obs. Date” entries listed in Table A1. | ccSee Sect. V.2 for a discussion of the source classifications. | |
| (h:m:s) | (d:m:s) | Phase | Bandpass/Flux | (m/m) | (yyyy-mm-dd) | () | (kpc) | ||
| 1 | G294.52-1.62 | 11:35:34.06 | -63:14:49.30 | J1123-6417 | J1107-4449 | 15.1/783.5 | 2022-06-02 | 1.40 | |
| 15.3/1210.6 | 2022-08-21 | ||||||||
| I11332-6258 | 11:35:32.23 | -63:14:46.80 | J1047-6217 | J1107-4449 | 15.1/1397.8 | 2024-05-31 | |||
| 2 | I12320-6122 | 12:34:53.38 | -61:39:46.90 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 4.17 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 3 | I12326-6245 | 12:35:34.81 | -63:02:32.10 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 4.21 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 4 | I12383-6128 | 12:41:17.32 | -61:44:38.60 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 4.12 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 5 | I12572-6316_1 | 13:00:24.03 | -63:32:31.90 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 11.63 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 6 | I12572-6316_2 | 13:00:28.73 | -63:32:37.30 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 11.63 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 7 | G305.21+0.21 | 13:11:13.78 | -62:34:41.90 | J1337-6509 | J1427-4206 | 15.1/783.5 | 2022-06-02 | 3.11 | |
| I13079-6218_1 | 13:11:13.73 | -62:34:40.20 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | |||
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 8 | I13079-6218_2 | 13:11:09.50 | -62:34:39.70 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 3.11 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 9 | I13080-6229 | 13:11:14.28 | -62:44:58.30 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 2.68 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 10 | I13111-6228 | 13:14:26.49 | -62:44:28.30 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 2.97 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 11 | I13134-6242 | 13:16:42.99 | -62:58:29.30 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 4.93 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 12 | I13140-6226 | 13:17:15.90 | -62:42:27.00 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 4.88 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 13 | I13291-6229_1 | 13:32:31.77 | -62:45:11.80 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 2.66 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 14 | I13291-6229_2 | 13:32:34.58 | -62:45:27.00 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 2.66 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 15 | I13291-6249 | 13:32:31.23 | -63:05:21.80 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 7.72 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 16 | I13295-6152 | 13:32:53.49 | -62:07:49.30 | J1337-6509 | J1617-5848 | 15.3/2516.8 | 2023-05-02 | 3.27 | |
| 15.1/1397.8 | 2024-06-02 | ||||||||
| 17 | G309.92+0.48 | 13:50:41.91 | -61:35:10.20 | J1337-6509 | J1427-4206 | 15.1/783.5 | 2022-06-02 | 5.17 | |
| I13471-6120 | 13:50:42.10 | -61:35:14.90 | J1408-5712 | J1617-5848 | 15.1/1397.8 | 2024-06-04 | |||
| 18 | G328.81+0.63 | 15:55:48.71 | -52:43:06.40 | J1603-4904 | J1617-5848 | 15.1/783.5 | 2022-05-30 | 2.56 | |
| 15.1/1301.6 | 2022-08-09 | ||||||||
| I15520-5234 | 15:55:48.39 | -52:43:09.80 | J1603-4904 | J1617-5848 | 15.1/1397.8 | 2024-06-02 | |||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 19 | G331.13-0.24 | 16:10:59.68 | -51:50:15.50 | J1603-4904 | J1617-5848 | 15.1/783.5 | 2022-05-30 | 5.29 | |
| 15.1/1301.6 | 2022-08-09 | ||||||||
| I16071-5142 | 16:10:59.01 | -51:50:21.60 | J1603-4904 | J1617-5848 | 15.1/1397.8 | 2024-06-02 | |||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 20 | G331.28-0.19 | 16:11:26.48 | -51:41:56.90 | J1603-4904 | J1617-5848 | 15.1/783.5 | 2022-05-30 | 5.31 | |
| 15.1/1301.6 | 2022-08-09 | ||||||||
| I16076-5134 | 16:11:27.2 | -51:41:56.90 | J1603-4904 | J1617-5848 | 15.1/1397.8 | 2024-06-02 | |||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 21 | G336.99-0.03 | 16:35:33.43 | -47:31:11.60 | J1631-4345 | J1617-5848 | 15.1/783.5 | 2022-05-30 | 7.95 | |
| I16318-4724 | 16:35:33.20 | -47:31:11.30 | J1650-5044 | J1617-5848 | 15.1/1397.8 | 2024-06-02 | |||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 22 | G339.88-1.26 | 16:52:04.83 | -46:08:34.40 | J1631-4345 | J1617-5848 | 15.1/783.5 | 2022-05-30 | 2.17 | |
| I16484-4603 | 16:52:03.99 | -46:08:24.60 | J1650-5044 | J1617-5848 | 15.1/1397.8 | 2024-06-02 | |||
| 15.1/1397.8 | 2024-06-03 | ||||||||
| 15.1/1397.8 | 2024-06-03 | ||||||||
Appendix B Image Alignment and Comparison of Difference Maps
| Source Name | Reference Epoch | Compared Epoch | (pix) | (pix) |
| I11332-6258 | 2022-06-02 | 2022-08-21 | +1.00 | +1.00 |
| 2022-06-02 | 2024-05-31 | 0.00 | 0.00 | |
| I12320-6122 | 2023-05-02 | 2024-06-02 | +0.28 | |
| I12326-6245 | 2023-05-02 | 2024-06-02 | +0.24 | +0.22 |
| I12383-6128 | 2023-05-02 | 2024-06-02 | +0.65 | |
| I12572-6316_1 | 2023-05-02 | 2024-06-02 | +0.26 | |
| I12572-6316_2 | 2023-05-02 | 2024-06-02 | +0.42 | |
| I13079-6218_1 | 2022-06-02 | 2023-05-02 | ||
| 2022-06-02 | 2024-06-02 | 0.00 | ||
| I13079-6218_2 | 2023-05-02 | 2024-06-02 | +0.01 | |
| I13080-6229 | 2023-05-02 | 2024-06-02 | +0.15 | +0.13 |
| I13111-6228 | 2023-05-02 | 2024-06-02 | +0.28 | |
| I13134-6242 | 2023-05-02 | 2024-06-02 | +0.00 | +0.12 |
| I13140-6226 | 2023-05-02 | 2024-06-02 | ||
| I13291-6229_1 | 2023-05-02 | 2024-06-02 | +0.28 | |
| I13291-6229_2 | 2023-05-02 | 2024-06-02 | +0.23 | |
| I13291-6249 | 2023-05-02 | 2024-06-02 | +0.07 | +0.02 |
| I13295-6152 | 2023-05-02 | 2024-06-02 | +0.54 | |
| I13471-6120 | 2022-06-02 | 2024-06-04 | +1.00 | 0.00 |
| I15520-5234 | 2022-05-30 | 2022-08-09 | +3.00 | 0.00 |
| 2022-05-30 | 2024-06-02 | +2.00 | +1.00 | |
| 2022-05-30 | 2024-06-03 | +3.00 | 0.00 | |
| I16071-5142 | 2022-05-30 | 2022-08-09 | 0.00 | +1.00 |
| 2022-05-30 | 2024-06-02 | +1.00 | ||
| 2022-05-30 | 2024-06-03 | +2.00 | ||
| I16076-5134 | 2022-05-30 | 2022-08-09 | +2.00 | 0.00 |
| 2022-05-30 | 2024-06-02 | +1.00 | +1.00 | |
| 2022-05-30 | 2024-06-03 | +1.00 | +1.00 | |
| I16318-4724 | 2022-05-30 | 2024-06-02 | 0.00 | |
| 2022-05-30 | 2024-06-03 | 0.00 | ||
| 2022-05-30 | 2024-06-03 | |||
| I16484-4603 | 2022-05-30 | 2024-06-02 | 0.00 | |
| 2022-05-30 | 2024-06-03 | 0.00 | ||
| 2022-05-30 | 2024-06-03 | 0.00 |
Note. — The first epoch is adopted as the reference epoch, and all spatial offsets are measured relative to this reference. The pixel scale of the images is 0.05 per pixel.
Appendix C Peak Intensities